OpenStudy (frostypaz1):

please show your work for this: x^2-8x+16=81

OpenStudy (lanhikari22):

Don't you know how to change form into (x+h)^2 ?

OpenStudy (hwyl):

you need to specify what is being asked for you to do with that equation

OpenStudy (frostypaz1):

it just says solve the equation

OpenStudy (hwyl):

solve for what?

OpenStudy (frostypaz1):

for x

OpenStudy (anonymous):

x^2-8x+16=81 x^2-8x+16-81=0 x^2-8x-65=0 (x-13)(x+5)=0 x=13,-5

OpenStudy (frostypaz1):

thank you!!!!